### v2

# iSeg: An Iterative Refinement-based Framework for Training-free Segmentation

Abstract
Stable diffusion has demonstrated strong image synthesis ability to given text descriptions, suggesting it to contain strong semantic clue for grouping objects. Inspired by this, researchers have explored employing stable diffusion for training-free segmentation. Most existing approaches either simply employ cross-attention map or refine it by self-attention map, to generate segmentation masks. We believe that iterative refinement with self-attention map would lead to better results. However, we empirically demonstrate that such a refinement is sub-optimal likely due to the self-attention map containing irrelevant global information which hampers accurately refining cross-attention map with multiple iterations. To address this, we propose an iterative refinement framework for training-free segmentation, named iSeg, having an entropy-reduced self-attention module which utilizes a gradient descent scheme to reduce the entropy of self-attention map, thereby suppressing the weak responses corresponding to irrelevant global information. Leveraging the entropy-reduced self-attention module, our iSeg stably improves refined cross-attention map with iterative refinement. Further, we design a category-enhanced cross-attention module to generate accurate cross-attention map, providing a better initial input for iterative refinement. Extensive experiments across different datasets and diverse segmentation tasks (weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation on synthetic dataset) reveal the merits of proposed contributions, leading to promising performance on diverse segmentation tasks. For unsupervised semantic segmentation on Cityscapes, our iSeg achieves an absolute gain of in terms of mIoU compared to the best existing training-free approach in literature. Moreover, our proposed iSeg can support segmentation with different kind of images and interactions. Our project page is publicly released at https://linsun449.github.io/iSeg/.
1 Introduction
Semantic segmentation [1, 2, 3] aims to differentiate the pixels into different semantic groups. In past decade, fully-supervised semantic segmentation has achieved significant progress where the models are trained on large-scale datasets with pixel-level annotations and fixed categories. This fully-supervised setting hinders zero-shot segmentation ability on unseen categories. Further, it is expensive to collect and annotate large-scale dataset with numerous categories [4] for model training. One straightforward solution is to develop a universal and training-free framework that exhibits strong generalization capabilities across arbitrary categories without any training on segmentation datasets.
Recent work of [5] introduces a vision-language model, CLIP, that learns the connection between text and images through training on large-scale image-text pairs, showing promising generalization capabilities in recognizing arbitrary categories. Inspired by this, several works [6, 7, 8, 9] explore employing the strong zero-shot ability of CLIP for semantic segmentation. Some works focus on directly utilizing the learned text-image mapping from pre-trained CLIP. For instance, OVSeg [6] first utilizes mask proposal generator to extract some class-agnostic masks, and then employs CLIP to classify cropped and masked images for open-vocabulary semantic segmentation. Some works employ text supervision similar to CLIP. For instance, GroupViT [8] introduces a hierarchical grouping vision transformer with text supervision, which is able to group different semantic regions without pixel-level annotations. These methods improve the generalization ability of semantic segmentation or reduce the reliance on pixel-level annotations in some degree. However, CLIP focuses on image-level information, limiting its capability in pixel-level tasks, especially under training-free setting.
Recently, researchers started to employ text-to-image stable diffusion model for segmentation. Stable diffusion [10] was initially designed to perform high-quality image synthesis. With the pre-training on large-scale text-image paired data, stable diffusion is able to generate high-quality images with reasonable structure according to given text descriptions. Inspired by this, researchers believed that stable diffusion contains rich semantic clue to group objects, and proposed various diffusion-based segmentation approaches. Some approaches [11, 12, 13, 14] treat stable diffusion as feature extractor, and re-train it on segmentation dataset. In contrast, other approaches [15, 16, 17] explore training-free framework by utilizing self-attention map and cross-attention map in stable diffusion. For instance, DiffSeg [15] iteratively merges self-attention maps at different positions and employs non-maximum suppression to assign the pixels into different objects. DiffuMask [18] generates initial object masks of synthetic images using cross-attention maps. DiffSegmenter [16] refines the cross-attention map with self-attention map for improved segmentation. These approaches present an initial attempt on employing self-attention and cross-attention maps for training-free segmentation tasks. However, we argue that it is still an open problem about how to fully harness self-attention and cross-attention maps for segmentation tasks.
To fully exploit the potential ability of cross-attention and self-attention maps in stable diffusion for training-free segmentation, we conduct a preliminary experiment, where we iteratively refine the cross-attention map with self-attention map as in Fig. 1. Given a text prompt (a) and an image (b), we employ stable diffusion to extract cross-attention map (c) of category bird. Subsequently, we perform a naive iteration strategy to refine the cross-attention map by iteratively multiplying self-attention map. After several iterations, we observe that the refined cross-attention map associated with category bird gradually has strong response on other regions, such as person and sandwich in (d). We argue that this degradation in refined cross-attention map is caused by self-attention map that aggregates global information from irrelevant regions after multiple iterations. To address this issue, we propose an iterative refinement framework by introducing an entropy-reduced self-attention (Ent-Self) module to impede the information aggregation from irrelevant regions. As in (e), our Ent-Self module is able to generate a stable and accurate cross-attention map of bird after multiple iterations. We further give a quantitative comparison of generated pseudo mask for weakly-supervised semantic segmentation in (f). Compared to naive iteration strategy, our Ent-Self module exhibits improved performance with the increment of iterations. In addition, we design a simple category-enhanced cross-attention (Cat-Cross) module to generate more accurate cross-attention map as the input fed to iterative refinement. The Cat-Cross module employs a weighting strategy to enhance the text embeddings of given category while suppressing other words. We perform the experiments on various datasets and segmentation tasks, which demonstrate method efficacy.
2 Related works
Semantic segmentation. Initially, researchers mainly investigated CNN-based networks. One of the pioneering approaches is FCN [19], which utilizes convolutional neural networks with skip-layer connections. Subsequently, numerous variants were developed, incorporating innovations such as spatial pyramid structures [1, 20], multi-layer feature fusion [21, 22, 23], and non-local enhancements [24, 25, 26]. With the success of vision transformer in image classification [27, 28], researchers began to explore transformer-based approaches. Some studies [29, 30, 3] focus on harnessing the powerful representational capabilities of transformer by employing it as backbone. Additionally, other researches [31, 32] conceptualize semantic segmentation as a set prediction task, utilizing the transformer as decoder.
Besides, researchers developed various semantic segmentation paradigms. Weakly-supervised semantic segmentation [33, 34, 35] solely relies on image-level category annotations to train models. For this task, many researchers employed CAM [36] to generate pseudo pixel-level annotations and then trained models with these annotations. Open-vocabulary semantic segmentation [37, 38, 8] performs semantic segmentation on unseen categories. In open-vocabulary semantic segmentation, researchers mainly explored to utilize CLIP [5] to segment unseen categories. Unsupervised segmentation [39, 40, 41] aims to group the pixels belonging to the same objects without using any annotations.
Diffusion models for segmentation. Diffusion models [42] have exhibited strong image synthesis ability across various text prompts, indicating their strong semantic clue for image segmentation. Inspired by this, researchers started to employ the pre-trained diffusion models for segmentation. For instance, VPD [13] views stable diffusion model as feature extractor, and fine-tunes them for semantic segmentation tasks. TADP [43] introduces image-aligned text prompts to condition diffusion model for segmentation. SLIME [44] adopts stable diffusion model with learnable text embeddings for few-shot segmentation. Additionally, some researchers [45] adopted diffusion framework for predicting accurate segmentation maps from noisy data.
Recently, some researchers explored training-free diffusion models for segmentation. One of the key idea is to employ self-attention map and cross-attention map from pre-trained diffusion model to generate masks of semantic categories. For instance, DiffuMask [18] first employs diffusion model to generate synthetic images, and calculates the averaged cross-attention map as the initial object masks. DiffSeg [15] explores grouping self-attention maps at different positions into distinct objects based on their similarities, followed by non-maximum suppression to assign pixels to individual objects. OVDiff [46] generates synthetic images by diffusion model, and uses a feature extractor to generate the foreground and background prototypes of different categories for open-vocabulary segmentation. Both Diffsegmenter [16] and T2M [17] refine cross-attention maps of different categories with self-attention maps. To improve cross-attention map, Diffsegmentor and T2M employ additional BLIP [47] and CLIP to carefully design text prompts, which are relatively complex in some degree. Though these aforementioned approaches exploit self-attention and cross-attention maps for training-free segmentation, we argue that these approaches do not fully utilize the semantic clue existing in self-attention and cross-attention maps. To address this issue, we propose an iterative training-free segmentation framework using pre-trained stable diffusion model.
3 Method
3.1 Preliminaries
Stable diffusion model. Text-to-image (T2I) diffusion model is a type of text-conditioned diffusion models [42], such as DALL-E [48], GLIDE [49], Imagen [50], and stable diffusion [10]. As one of representative T2I models, stable diffusion generates high-quality images with text prompts, which is trained on large-scale text-image paired dataset (e.g., LAION-5B [51]). It comprises three components: (1) A pre-trained VAE [52] encoder and decoder for image compression and restoration. (2) A text encoder that encodes text descriptions into latent semantic space. (3) A transformer-based denoising UNet [53] that recovers images from noisy data according to text descriptions.
Attention modules. Denoising UNet in stable diffusion integrates self-attention and cross-attention mechanisms. The self-attention models the relation between different spatial positions of visual features, while the cross-attention models the relation between visual features and text embeddings. The attention map in both self-attention and cross-attention operations can be calculated by the query and the key as , where is normalization factor. Assuming that the visual features has the spatial size of , and the text embeddings has the length of . In self-attention operation, the query and key are both generated by visual features via linear layers. The generated self-attention map is denoted as , which can reflect spatial similarities of different positions within image. In cross-attention operation, the query and key are respectively generated by visual features and text embeddings via linear layers. The generated cross-attention map is denoted as , capturing the similarities between image and text.
To achieve training-free segmentation, a straightforward way is setting the text prompts based on given category names, such as ‘A photograph of cat and dog’, then performing the forward process of stable diffusion and finally generating the corresponding cross-attention map between visual features and text embeddings. Afterwards, we can predict segmentation map based on the obtained cross-attention map associated with corresponding category name. Instead of directly using cross-attention map, some researchers [44] proposed to generate refined cross-attention map with self-attention map, which can be written as
Compared to original cross-attention map, the refined cross-attention map incorporates spatial similarities between different positions existing in self-attention map [44], which is useful to perform accurate segmentation.
3.2 Motivation
Problem. As mentioned above, self-attention map can be used to refine the cross-attention map. Therefore, a natural question is that, can we improve cross-attention map iteratively with self-attention map, formulated as , where is the index of iterations. We initially believe that this iterative refinement would yield improved segmentation performance. However, as in Fig. 1(f), it presents performance degradation after several iterations. To better analyze this question, we perform the experiment in Fig. 2. Given the input image (a) and cross-attention map (b), we present the refined cross-attention map at different iterations in (c). It can be observed that the refinement through multiple iterations can not guarantee to generate better refined cross-attention map. We divide the refined results into three situations. The first situation is that the cross-attention map becomes better with the increasing iterations, such as cross-attention map of dog in first row. The second one is the cross-attention map becomes larger, such as the cross-attention map of sheep in the second row. The last one is that the cross-attention map shifts to other regions, exemplified by the cross-attention map of dog showing strong response on the person in third row.
Analysis. Ideally, we argue the refinement with multiple iterations at least presents a stable cross-attention map around corresponding category if the self-attention map is relatively accurate. We further perform a simple experiment by refining cross-attention map with ground-truth self-attention map in Fig. 2(d). The ground-truth self-attention map is generated by self-correlation of ground-truth masks, which is binary and exclusively has responses within object regions. By using ground-truth self-attention map, the refined cross-attention map becomes stable and accurate after only one iteration. This observation demonstrates that an accurate self-attention map is crucial for accurately refining cross-attention map. As shown in (e), we compare the predicted self-attention map (middle) and ground-truth self-attention map (right) at a given point . Compared to ground-truth self-attention map, predicted self-attention map is relatively noisy and has some weak responses at irrelevant regions. As a result, the refinement with predicted self-attention map aggregates irrelevant global information into the refined cross-attention map. Therefore, we argue that it is important to reduce irrelevant global information within self-attention map for iterative refinement.
3.3 Our iterative refinement framework
Here we introduce our proposed iterative refinement framework for training-free segmentation, named iSeg. We employ pre-trained stable diffusion model for feature extraction, and perform segmentation using the generated self-attention and cross-attention maps with iterative refinement. To address performance degradation issue in iterative refinement, we introduce an entropy-reduced self-attention (Ent-Self) module to suppress irrelevant global information within self-attention map. In addition, we design a category-enhanced cross-attention (Cat-Cross) module to generate more accurate cross-attention map as the input of iterative refinement. The algorithm is detailed in Appendix A.
Iterative refinement. Fig. 3 shows the overall architecture of our proposed iSeg. Given an image , we feed it into a VAE encoder to extract visual latent features . At the same time, we employ a text encoder to generate text embeddings from corresponding input texts. Afterwards, we feed the visual latent features and text embeddings into the denoising UNet, which is formulated as
where represents the denoising U-net, and indicates the timestep that is set as a fixed value like [16]. Then, we can extract self-attention and cross-attention maps at different layers of denoising UNet. Similar to [44], we only select the high-resolution self-attention map at last decoder layer, and upsample and fuse the cross-attention maps at different decoder layers as initial cross-attention map . For self-attention map , we introduce an entropy-reduced self-attention (Ent-Self) module to reduce irrelevant global information within original self-attention map, where the entropy-reduced self-attention map is represent as . Then, we perform the iterative refinement using entropy-reduced self-attention map as
where represents the total number of iterations, and is the original cross-attention map . After each iteration, we add a normalization operation. For simplicity, we do not show the normalization in Eq. 3 and Fig. 3. Finally, we can extract the segmentation map of a category at the index of text prompts as . It is worth noting that an accurate initial cross-attention map is also important, which provides a better input for iterative refinement. Therefore, instead of original cross-attention map, we design a simple category-enhanced cross-attention (Cat-Cross) module to obtain more accurate cross-attention map fed to iterative refinement.
Entropy-reduced self-attention module. As mentioned above, the weak responses of irrelevant regions within self-attention map contain the global information and interfere with the refined cross-attention map. To address this issue, we introduce an entropy-reduced self-attention (Ent-Self) module. Following [54], we calculate self-attention entropy by viewing the attention distribution as a probability mass function of a discrete random variable, which can be written as
The self-attention entropy has the maximum value when for all elements, and has the minimum value when is binary like ground-truth self-attention map mentioned above. This means that the higher the self-attention entropy is, the more global information will be, and thereby the more likely irrelevant information will be. To reduce the irrelevant information, for a given element in self-attention map, the entropy gradient can be calculated as
Finally, we calculate the entropy-reduced self-attention map by updating the original self-attention map along negative gradient direction as
where represents the updating factor. The entropy-reduced operation makes the value of near to one becomes larger, and makes the value of near to zero becomes smaller. As a result, the entropy-reduced operation can suppress the weak responses within self-attention map corresponding to irrelevant global information, and does not reduce the strong responses corresponding to the local region.
Category-enhanced cross-attention module. The accurate cross-attention map presents a good initial state for iterative refinement. Ideally, if we have a perfect cross-attention map, we do not require self-attention refinement. As show in Fig. 4(b), the original cross-attention maps usually have the strong responses on irrelevant regions, such as other categories or background. To improve the quality of cross-attention map fed to the refinement, we introduce a simple category-enhanced cross-attention (Cat-Cross) module, which makes the attention focus more on the text features belonging to the given categories by a factor of . Specifically, for a set of given categories for segmentation, we generate a weighting vector as
where is larger than . With the weighting factor , we generate weighted cross-attention map as
where means element-wise multiplication, the query represents visual features, and the key represents the text embeddings.
4 Experiments
We perform the experiments on weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation for synthetic dataset. Following the prior works [15, 16], we mainly use pixel accuracy (ACC) and mean intersection over union (mIoU) to evaluate segmentation performance.
| Method | Publication | Training | VOC | COCO | |
| CNN-based | IRN [55] | CVPR2019 | 66.5 | 42.4 | |
| AdvCAM [56] | CVPR2021 | 55.6 | 35.8 | ||
| BAS [57] | IJCV2023 | 57.7 | 36.9 | ||
| HSC [58] | IJCAI2023 | 71.8 | - | ||
| Transformer-based | MCTformer [59] | CVPR2022 | 61.7 | - | |
| MCTformer+ [60] | arXiv2023 | 68.8 | - | ||
| ToCo [61] | CVPR2023 | 72.2 | - | ||
| WeakTr [62] | arXiv2023 | 66.2 | - | ||
| CLIP-based | CLIMS [9] | CVPR2022 | 56.6 | - | |
| CLIP-ES [7] | CVPR2023 | 70.8 | 39.7 | ||
| DiffSegmenter [16] | arXiv2023 | 70.5 | - | ||
| T2M [17] | arXiv2023 | 72.7 | 43.7 | ||
| baseline (Ours) | - | 66.0 | 39.4 | ||
| Diffusion-based | iSeg (Ours) | - | 73.6 | 44.7 |
Implementation details. We employ the pre-trained stable diffusion V2.1 as the base model, and feed the input images into the pre-trained model for segmentation without any training on segmentation dataset. The input images are resized to 512512 pixels. All segmentation tasks are evaluated on a single NVIDIA RTX 3090 GPU with 24G memory. During inference, only 5G memory is required when using half precision with a batch size of 1, resulting in an approximate inference time of 0.16 seconds per image. For weakly-supervised semantic segmentation, we evaluate pseudo mask generation and use the categories of corresponding training images to generate text prompts. For open-vocabulary semantic segmentation, we employ pre-trained CLIP to generate text prompts. For unsupervised segmentation and mask generation for synthetic dataset, we employ our entropy-reduced self-attention module to refine the segmentation results of DiffSeg [15] and DiffuMask [18]. Please refer to Appendix B for more implementation details of different segmentation tasks.
| Method | Publication | Training | VOC | Context | Object | |
| CLIP-based | ReCo [41] | NeurIPS2022 | 25.1 | 19.9 | 15.7 | |
| MaskCLIP [63] | ECCV2022 | 38.8 | 23.6 | 20.6 | ||
| SegCLIP [64] | ICML2023 | 52.6 | 24.7 | 26.5 | ||
| CLIPpy [65] | ICCV2023 | 52.2 | - | 32.0 | ||
| ViewCo [66] | ICLR2023 | 52.4 | 23.0 | 23.5 | ||
| TagCLIP [67] | AAAI2024 | 64.8 | - | - | ||
| OVDiff [46] | arXiv2023 | 67.1 | 30.1 | 34.8 | ||
| DiffSegmenter [16] | arXiv2023 | 60.1 | 27.5 | 37.9 | ||
| baseline (Ours) | - | 61.4 | 26.4 | 36.1 | ||
| Diffusion-based | iSeg (Ours) | - | 66.0 | 30.6 | 38.2 |
4.1 State-of-the-art comparison
Weakly supervised semantic segmentation. Table 1 compares our training-free method iSeg and some weakly-supervised methods, including CNN-based, Transformer-based, CLIP-based, and Diffusion-based approaches, on both PASCAL VOC [68] and MS COCO [69] training sets. Here, we only compare the quality of generated pseudo masks. Our proposed iSeg achieves the mIoU scores of 73.6% and 44.7% on PASCAL VOC and MS COCO, respectively. Compared to diffusion-based training-free DiffSegmenter [16] and T2M [17], our iSeg has 3.1% and 0.9% improvements in terms of mIoU on PASCAL VOC. Our proposed iSeg also outperforms the training-based approaches. For instance, compared to ToCo [61] and WeakTr [62], our proposed iSeg has 1.4% and 7.4% improvements in terms of mIoU on PASCAL VOC.
Open-vocabulary semantic segmentation. Table 2 compares our proposed method with some open-vocabulary semantic segmentation methods, including CLIP-based and diffusion-based approaches, on PASCAL VOC validation set, PASCAL Context validation set [70], and MS COCO Object validation set. Our proposed iSeg outperforms the training-free DiffSegmenter [16] by an absolute gain of , , and on these three datasets.
| Method | Training | Cityscapes | COCO-Stuff-27 | ||
| ACC | mIoU | ACC | mIoU | ||
| IIC [39] | 47.9 | 6.4 | 21.8 | 6.7 | |
| MDC [71] | 40.7 | 7.1 | 32.3 | 9.8 | |
| PICLE [72] | 65.5 | 12.3 | 48.1 | 13.8 | |
| STEGO [40] | 73.2 | 21.0 | 56.9 | 28.2 | |
| MaskCLIP [63] | 35.9 | 10.0 | 32.2 | 19.6 | |
| RoCo [41] | 74.6 | 19.3 | 46.1 | 26.3 | |
| DiffSeg [15] | 76.0 | 21.2 | 72.5 | 43.6 | |
| iSeg (Ours) | 78.7 | 25.0 | 74.5 | 45.2 | |
Unsupervised segmentation. Table 3 compares our method with some unsupervised segmentation methods on Cityscapes [73] validation set and COCO-Stuff-27 [74] validation set. We employ our iterative refinement as a post-processing, and integrate it into unsupervised DiffSeg [15] to improve segmentation results. Our proposed iSeg can also present stable improvement on both Cityscapes and COCO-Stuff-27. For instance, our iSeg outperforms DiffSeg by a large gain of in terms of ACC and in terms of mIoU on Cityscapes.
Mask generation on synthetic dataset. Fig. 5 presents some qualitative results of our proposed iSeg and DiffuMask [18]. The synthetic images are from DiffuMask [18]. DiffuMask first generates coarse pseudo masks using cross-attention map, and then train an AffinityNet to generate more precise pseudo masks. Compared to DiffuMask, our proposed iSeg has more accurate segmentation results without using any training processing. More results on synthetic images can be found in appendix C.
4.2 Ablation study
| Ent-Self | Cat-Cross | mIoU |
| 66.0 | ||
| 70.1 | ||
| 68.9 | ||
| 73.6 |
Impact of integrating two proposed modules. Table LABEL:ablation:modules shows the results on PASCAL VOC training set. When integrating entropy-reduced self-attention (Ent-Self) module, it achieves a performance improvement of 4.1%. When integrating category-enhanced cross-attention (Cat-Cross) module, it presents a performance improvement of 2.9%. By integrating both Ent-Self and Cat-Cross into the baseline, it totally improves the baseline by 7.6%, which demonstrates that two proposed modules are complementary. Specifically, Cat-Cross module provides a better input for refinement, while Ent-Self module avoids the performance degradation during refinement.
Hyper-parameter settings. There are some hyper-parameters in our method, including the updating factor in our Ent-Self module, the number of iterations , and the weighting factor in our Cat-Cross module. Fig. 6(a) presents the impact of and . When (namely using original self-attention map), iterative refinement presents a fast performance drop after several iterations. With a relatively large (e.g., 0.05), iterative refinement initially improves performance, and then keeps relatively stable performance. When and , it achieves best performance. Fig. 6(b) presents the impact of weighting factor . Increasing initially leads to improved performance but eventually results in performance drop. The best performance is achieved when .
4.3 Segmentation with different kinds of images and interactions
Fig. 7 shows that our method supports segmentation on different kinds of images and interactions. At left, our method performs accurate segmentation of given categories on natural, synthetic, clipart, painting and sketch images. At right, our method can support different interactions to perform segmentation, where the initial segmentation map is generated by unsupervised segmentation maps, point, box, and line, and then fed to our iterative refinement with Self-Ent module. More results are in Appendix C. We also provide a video demo in supplementary material.
5 Conclusion
In this paper, we propose an iterative refinement framework for training-free segmentation using pre-trained stable diffusion model, named iSeg. Our proposed iSeg introduces two novel modules: entropy-reduced self-attention module and category-enhanced cross-attention module. The entropy-reduced self-attention module aims to suppress irrelevant global information within self-attention map, while the category-enhanced cross-attention module pays more attention to the features of given segmented categories. With these two proposed modules, our proposed method can generate accurate segmentation masks without any training on segmentation datasets. We perform the experiments on various datasets and segmentation tasks to demonstrate the efficacy of our proposed method.
Limitations and future work. Currently, we set the number of iterations in our iterative refinement as a fixed number, which is not optimal. In fact, the number of iterations can be dynamically changed according to the image. In future, we will explore image-aware iterative refinement framework.
Appendix A Algorithms
Algorithm 1 presents the pseudo code about how to refine the cross-attention by entropy-reduced self-attention iteratively using our proposed iSeg.
Appendix B More implementation details
Here we present more implementation details of our proposed method on different segmentation tasks, including weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation for synthetic dataset.
Weakly-supervised semantic segmentation. In weakly-supervised semantic segmentation, one of the key is to generate pseudo pixel-level masks for training set. We compare our training-free method with existing approaches for pixel-level mask generation on two datasets, including PASCAL VOC 2012 [68] and MS COCO 2014 [69]. Specifically, we use the categories existing in image to generate a template such as ‘A photograph of [CLS A] and [CLS B] and other objects and background’. Here we compose the sentence with some general words, such as ‘object’ and ‘background’ rather than specific background category associated with the specific dataset similar to CLIPES [7].
Open-vocabulary semantic segmentation. Open-vocabulary semantic segmentation aims to perform segmentation of arbitrary categories. We compare our training-free method with existing approaches on three datasets, including PASCAL VOC 2012 (VOC), PASCAL-Context (Context), and MS COCO Object (Object). To perform open-vocabulary semantic segmentation, we employ the pre-trained CLIP to generate image-level labels as text prompts following TagCLIP [67]. Afterwards, we keep the same settings as weakly-supervised semantic segmentation task to obtain the final mask.
Unsupervised segmentation. We compare our training-free method with other unsupervised segmentation approaches on two datasets, including COCO-Stuff-27 [74] and Cityscapes [73]. We integrate our proposed method into the training-free approach DiffSeg [15] as a post-processing. Specifically, we treat the segmentation results of DiffSeg as initial cross-attention maps, and employ our proposed entropy-reduced self-attention module to perform iterative refinement, while the other settings are consistent with DiffSeg.
Mask generation for synthetic dataset. Recently, some researchers have explored generating synthetic dataset for semantic segmentation. In this task, it is important to generate the mask labels for synthetic images. DiffuMask [18] utilizes cross-attention maps in diffusion model to generate pseudo masks for synthetic images, which are further refined by AffinityNet. Here, we can directly apply our iterative refinement to improve the initial pseudo masks generated by DiffuMask on synthetic dataset without requiring any additional refinements like AffinityNet.
Appendix C Qualitative results
Cross-attention map after different iterations. Fig. 8 visualizes some refined cross-attention maps of our proposed iSeg at different refinement iterations. With the increment of iterations ranging from 0 to 8, our proposed iSeg can generate more accurate cross-attention maps of corresponding objects.
Results on synthetic images. Fig. 9 presents some segmentation results of our proposed iSeg on synthetic images generated by DiffuMask [18]. It can be observed that our proposed method can generate accurate segmentation masks for diverse synthetic images.
Results on natural images. Fig. 10 presents some segmentation results of our proposed iSeg on PASCAL VOC 2012 and MS COCO. It can be observed that our proposed method can generate accurate segmentation masks for various categories.
Results of unsupervised segmentation. Fig. 11 presents some unsupervised segmentation results of our proposed iSeg on COCO-Stuff-27 datasets. Our proposed method is able to accurately group the pixels into different objects.
Cross-domain results. Fig. 12 presents some kind of image segmentation on DomainNet [75]. Our proposed method is able to accurately segment different kinds of objects including sketch, painting and clipart according to given category.
Results using different interactions. Fig. 13 presents more segmentation results using different interactions, such as points, lines, and boxes.